### v4

# Exploring the Impact: How Decentralized Exchange Designs Shape Traders’ Behavior on Perpetual Future Contracts

Abstract
In this paper, we analyze traders’ behavior within both centralized exchanges (CEXs) and decentralized exchanges (DEXs), focusing on the volatility of Bitcoin prices and the trading activity of investors engaged in perpetual future contracts. We categorize the architecture of perpetual future exchanges into three distinct models, each exhibiting unique patterns of trader behavior in relation to trading volume, open interest, liquidation, and leverage. Our detailed examination of DEXs, especially those utilizing the Virtual Automated Market Making (VAMM) Model, uncovers a differential impact of open interest on long versus short positions. In exchanges which operate under the Oracle Pricing Model, we find that traders primarily act as price takers, with their trading actions reflecting direct responses to price movements of the underlying assets. Furthermore, our research highlights a significant propensity among less informed traders to overreact to positive news, as demonstrated by an increase in long positions. This study contributes to the understanding of market dynamics in digital asset exchanges, offering insights into the behavioral finance for future innovation of decentralized finance.
keywords:
Blockchain. Perpetual Futures. Volatility. DeFi.
1 Introduction
Over the past decade, blockchain technology has seen rapid development, attracting over 200 million users globally. Among its applications, cryptocurrency derivative trading, particularly perpetual futures, has become prominent, with daily trading volumes exceeding \$100 billion. Perpetual futures, introduced into the cryptocurrency market by BitMEX (https://bitmex.com/) in 2016 and conceptualized by Robert J. Shiller[19], represent a novel variant of traditional futures contracts. Unlike their traditional counterparts, perpetual futures do not have a fixed delivery date with their prices pegged to spot prices of the underlying assets, allowing for indefinite trading without the need for contract rollovers.Investors show a preference for perpetual futures over traditional futures due to their efficiency in hedging and speculation, high leverage, and the absence of delivery and rollover requirements. Before 2022, investors trade perpetual futures on centralized exchanges (CEXs), but following the collapse of FTX there emerged many decentralized exchanges (DEXs) carrying perpetual futures, which are operated with more elements on-chain. Since the DEXs are constructed with smart contracts involved, more transparency can be ensured, helping guard against the fraudulence incurred in CEXs.
Although both running perpetual futures, CEXs and DEXs are considerably different in terms of their order matching and price discovering, leading to different traders’ behavior and implications on the market micro-structure. The design of the current CEXs and DEXs can be classified into three exchange models according to the way orders are matched: Limit-Order-Book (LOB) Model, Virtual Market Making (VAMM) Model, and Oracle Pricing Model. Unlike the case of LOB Model, which adopt the limit-order book to form prices, orders submitted by traders in exchanges of VAMM Model are matched with liquidity providers  [^1] Liquidity providers deposit assets in liquidity pools facilitated by smart contracts, where traders execute transactions. and the prices are formed based on the relative abundance of assets in the liquidity pool. While in exchanges of LOB and VAMM Model traders’ trading activity help form the price, traders in exchanges of Oracle Pricing Model accept the price offered by the Oracle and thus act as pure price takers. Although liquidity providers are still passively act as the trading counterparty, the pricing is governed by a Price Oracle, which assimilates prices from various spot exchanges, thereby rendering transactions in exchanges of Oracle Pricing Model incapable of impacting the price directly [^2] Details about the design of different exchange models can be found in Appendix A or the Systems of Knowledge paper [8].. Therefore, in this paper we ask how those differences in exchanges of difference exchange models affect price formation and investors’ behavior and what kinds of empirical implications can be provided to the potential theoretical models for those exchange models.
We investigate the distinct behaviors of traders on exchanges within different models by examining the relationship between the price volatility of the underlying asset, i.e., Bitcoin, and various trading activities, i.e., trading volume, open interest, liquidation, and leverage. Our study encompasses prominent exchanges of LOB Model such as Binance (https://www.binance.com/en) and Bybit (https://www.bybit.com), alongside exchanges of Oracle Pricing Model like GMX (https://gmx.io) and GNS (https://gains.trade), and Perpetual Protocol V2 (https://perp.com) which employs VAMM Model. The empirical evidence shows that, for exchanges of LOB Model (Binance and Bybit), price volatility is positively related with trading volumes but negatively related with open interests, consistent with findings in the traditional future markets. While the effects of trading activity on price volatility in exchanges of LOB Model reflect the market depth and can be explained by Kyle’s model [13], the VAMM Model (Perpetual Protocol V2) introduces a nuanced dynamic in the impact of open interest, varying between long and short positions. This asymmetry is attributable to the VAMM’s price formation mechanism, where the rate of change in an asset’s relative price inversely correlates with its abundance in the liquidity pool. Consequently, market depth increases with rising open interests in short positions, as the underlying asset accumulates in the liquidity pool, and the reverse holds true for long positions. While in exchanges of LOB and VAMM Model traders’ trading activity help form the price, traders in exchanges of Oracle Pricing Model (GMX and GNS) accept the price offered by the Oracle and thus act as pure price takers. Therefore, trading activity should be interpreted as purely traders’ reaction to the price change of the underlying asset. When the price become more volatile, trading volumes increase while open interests decrease, with change in long and short positions different. These empirical estimations can be explained by the predictions based on Shalen’s dispersion of beliefs model (1993) [18], which address the asymmetry of information traders can access. We also find that uninformed traders tend to overreact more to positive news than negative, evidenced by more of the long positions accumulated.
Existing literature evaluates cryptocurrency future contracts mainly in terms of their relationship with the spot market. Akyildirim et al. [1] studied the impact of Bitcoin futures on the cryptocurrency market, especially the introduction of CME and CBOE futures contracts in December 2017. Alexander et al. [2] found that BitMEX derivatives lead the price discovery process over Bitcoin spot markets. Hung et al. [12] identifies substantial pricing effects and breakpoints in market efficiency, indicating the dominant role of Bitcoin futures in price discovery compared to spot markets. As a special case of future contracts, perpetual future contracts are scarcely investigated, although with much higher trading volume. Besides the theoretical discussion on the arbitrage between perpetual future markets and spot markets [11], there lacks enough empirical works examining perpetual future traders’ behavior, especially in DEXs. After the pioneering work by Soska et al. [20], which conducted the first analysis on the trader profile for perpetual future contracts in BitMEX, i.e., a CEX, Alexander et al. [4] constructed the optimal hedging strategy with empirical corroboration.
Since some DEXs of perpetual futures adopt VAMM Model, which utilize Automated Market Making (AMM) as the core, our study also shed lights on the effect of traders specifically derived by AMM. There is a emerging sort of research on the economic implication of AMM, most of which, however, focus on the incentive of liquidity provision, driven by the relationship between transaction fee and impermanent loss [16, 10, 14, 6]. While the limited existing empirical studies are based on the spot market [14, 9], i.e., Uniswap, our study contribute to the understanding of AMM by conducting the first analysis on the future market supported by AMM from the perspective of traders, enlightening the effect driven by AMM in a new context.
While the primary focus of our research is on Decentralized Exchanges (DEXs) operating with perpetual future contracts, the findings and theoretical frameworks developed herein have broader applicability, extending to other trading paradigms within DEXs that exhibit the fundamental bifurcation between Traders and Liquidity Providers. This is exemplified in platforms such as PancakeSwap (https://pancakeswap.finance) and Uniswap (https://uniswap.org). Furthermore, as the Decentralized Finance (DeFi) ecosystem evolves, the proliferation of synthetic assets is anticipated to enhance the diversification of portfolios managed on DEXs. The insights gleaned from this study are thus of paramount importance to both academic researchers and industry practitioners.
2 Data and Methodology
GMX constructs its trading system on Arbitrum and Avalanche  [^3] Arbitrum and Avalanche are two chains to address scalability of Ethereum (https://vitalik.eth.limo/general/2023/10/31/l2types.html), while detailed explanation can be found at: arbitrum.io and www.avax.network., and we focus on tradings on Arbitrum from 31 August 2021 to 26 September 2023, the data of which are retrieved from Dune (https://dune.com/queries/3110823). Similarly, GNS is based on Arbitrum and Polygon  [^4] Information about Polygon can be found at: polygon.technology., and we choose the perpetual futures on Polygon from 20 December 2021 to 26 September 2023 to examine, the data of which are retrieved by our constructed queries on Dune (https://dune.com/queries/2829105 and https://dune.com/queries/2825251). We also retrieve data for Perpetual Protocol V2 (based on Etherum mainnet) from 27 November 2021 to 26 September 2023 by constructing the query on Dune (https://dune.com/queries/3039245).
We retrieve data on Binance and Bybit at the daily-level frequency using API provided by Coinalyze (https://coinalyze.net/) from 05 August 2020. Data of BTC pricing at the daily-level are collected from CryptoCompare with API (https://cryptocompare.com/).
It is paramount to recognize that the pricing mechanisms governing perpetual futures contracts are intentionally designed to approximate the spot price of the underlying asset. Consequently, traders engaged in the trading of perpetual futures are confronted with prices inherently tethered to the underlying asset’s market valuation[11]. For the purpose of our analysis, and without loss of generality, we direct our focus toward an examination of how traders operating within the perpetual futures market respond to variations in the price of the underlying asset, specifically Bitcoin.
To establish an estimation for the daily volatility of Bitcoin’s price, this study employs the extreme-value volatility estimator, proposed by Garman and Klass (1980) [21], as follows:
$\begin{split}\hat{\sigma}_{t}=\left\{0.5\times\left(\ln\left(P_{t,H}/P_{t,L}\right)\right)^{2}\right.-\left.(2\ln(2)-1)\left(\ln\left(P_{t,O}/P_{t,C}\right)\right)^{2}\right\}^{1/2},\end{split}$ (1)
where $P_{t,H},P_{t,L},P_{t,O}$, and $P_{t,C}$ are the high, low, opening, and closing prices of Bitcoin on date $t$, respectively.
Previous research has consistently shown that variables related to traders’ activities, such as trading volumes and open interests, exhibit significant correlation [5], suggesting their predictability based on the previous values. Consequently, these variables are often analyzed by dividing them into expected and unexpected components. The expected component encompasses information about current trends, while the unexpected component captures unforeseen changes in traders’ behavior. Hedgers, who adjust positions infrequently with a long-term focus, exhibit predictable behavior captured in the expected component. In contrast, speculators, who frequently adjust positions to manage risk, are represented by the unpredictable, unexpected component of activity. Following the methodology of Bessembinder and Seguin (1993) [5], we partition trading volume, open interest, volume of liquidated positions, and daily average leverage using ARIMA$(p,k,q)$ model, where fitted values are regarded as the expected component and its residuals as the unexpected component.
To explore how investors’ behavior vary with volatility, we adapt the model proposed by Wang and Yau (2002) [24] to include daily volumes of liquidated short and long positions separately:
$\hat{\sigma}_{t}=$ $\mu+\sum_{i=1}^{m}\phi_{i}\hat{\sigma}_{t-i}+\sum_{j=1}^{3}\alpha_{j}EA_{j,t}+\sum_{j=1}^{3}\beta_{j}UA_{j,t}+\sum_{k=1}^{2}\gamma_{k}EL_{k,t}+$ (2)
$\sum_{k=1}^{2}\lambda_{j}UL_{k,t}+\varepsilon_{t},$
where $\hat{\sigma}_{t}$ is the estimated volatility on day $t$. $EA_{j,t}$ and $UA_{j,t}$ denote the expected and unexpected trading activity respectively, with $j$ represents different trading activity when assigned different values: $j=1$ for open interest on long positions, $j=2$ for open interest on short positions, and $j=3$ for trading volumes. $EL_{k,t}$ and $UL_{k,t}$ denote represent the expected and unexpected daily liquidated volumes for long ($k=1$) and short ($k=2$) positions. The model accounts for volatility persistence through lagged volatility estimates, with the lag structure ($m$) determined by the Akaike Information Criterion. Given the established correlation between trading activities and volatility [24], these variables are included as controls. This analysis using Eq. (2) is applied to perpetual futures from all three DEXs and two CEXs, given the availability of trading volume, open interest, and liquidation data.
Anticipating that traders are inclined to modify their leverage strategies in response to heightened liquidation risks under increased price volatility of the underlying asset, this study incorporates an analysis of leverage adjustments. Research focusing on centralized exchanges (CEXs) often omits leverage considerations due to the general unavailability of such data [22, 20]. However, leveraging transaction-level data from GMX and GNS, this investigation calculates and integrates the daily average leverages for both long and short positions into the analytical framework. This is achieved by adapting the existing volatility estimation model, as delineated in Equation (2):
$\hat{\sigma}_{t}=$ $\mu+\sum_{i=1}^{m}\phi_{i}\hat{\sigma}_{t-i}+\sum_{j=1}^{3}\alpha_{j}EA_{j,t}+\sum_{j=1}^{3}\beta_{j}UA_{j,t}+\sum_{k=1}^{2}\gamma_{k}EL_{k,t}+$ (3)
$\sum_{j=k}^{2}\lambda_{j}UL_{k,t}+\sum_{k=1}^{2}\rho_{k}ELV_{k,t}+\sum_{k=1}^{2}\omega_{k}ULV_{k,t}+\varepsilon_{t},$
where $ELV_{k,t}$ and $ULV_{k,t}$ denote expected and unexpected average leverage for long ($k=1$) and short positions ($k=2$), respectively, on day $t$. Notably, due to the cross-margining feature in Perpetual Protocol v2, where a collective pool of funds backs each position  [^5] Documents of Leverage and Margin Ratio for Perpetual Protocol v2: support.perp.com/hc/en-us/articles/5257393633945-Leverage-Margin-Ratio., the specific leverage level for each position remains undefined. Consequently, the leverage analysis using Equation (3) is confined to perpetual futures from GMX and GNS.
3 Empirical Results
3.1 Empirical results and analysis for LOB Model (Binance and Bybit)
Table 1: Regression Results on Eq. (2)
[…]
Binance Bybit [^8] Since for CEXs each future contract is matched with a buyer and a seller, resulting in the same dollar value of open interests for long and short, we regress price volatility only on open interest on short. Perpetual Protocol V2 GMX GNS
Intercept [^9] In each cell, the t-statistics are in the parentheses. *, **∗ ∗, and ***∗ ∗ ∗ denote significance at 0.1, 0.05, and 0.01 level, respectively. 0.011 (6.27)*** 2.145 (0.15) 0.009 (3.48)*** 1.438e-12 (16.82)*** 0.017 (8.82)***
Lagged volatility 0.200 (10.16)*** 0.210 (17.19)*** 0.252 (8.19)*** 0.256 (8.10)*** 0.371 (10.57)***
        Trading Activity:
Expected trading volume 3.06e-08 (12.87)*** 6.041e-08 (8.91)*** 2.59e-09 (8.76)*** 1.186e-10 (6.30)*** 1.222e-11 (0.19)
Unexpected trading volume 5.943e-08 (24.07)*** 1.409e-07 (15.62)*** 3.62e-09 (14.42)*** 1.29e-10 (9.83)*** 4.53e-10 (6.65)***
Expected OI on short -1.11e-12 (-2.24)*** -8.55e-12 (-8.69)*** -1.678e-09 (-4.40)*** -4.765e-10 (-7.36)*** -2.059e-10 (-1.99)**
Unexpected OI on short -2.365e-11 (-8.93)*** -3.138e-11 (-4.65)*** 1.125e-08 (4.53)*** -2.808e-10 (-3.27)*** -4.654e-10 (-3.08)***
Expected OI on long - - 3.569e-09 (3.98)*** -1.88e-10 (-5.99)*** 2.98e-11 (0.44)
Unexpected OI on long - - 1.481e-09 (0.161) -1.75e-10 (-2.16)** -4.261e-11 (-0.29)
        Liquidation:
Expected liquidated volume on short positions 2.864e-10 (5.46)*** -2.832e-10 (-2.32)** -5.983e-07 (-2.09)** 2.449e-09 (1.09) -1.483e-09 (-1.26)
Unexpected liquidated volume on short positions 1.655e-10 (6.95)*** 5.098e-10 (7.22)*** -3.032e-08 (-1.61) 2.007e-09 (1.61) 3.386e-10 (0.835)
Expected liquidated volume on long positions 8.534e-11 (3.00)*** 1.672e-09 (11.96)** 5.008e-07 (4.16)*** 1.82e-07 (17.71)*** 7.295e-09 (3.26)***
Unexpected liquidated volume on long positions 7.143e-11 (6.82)*** 1.278e-09 (16.62)*** 1.482e-07 (5.39)*** 2.593e-09 (5.08)*** 1.448e-09 (1.46)
Adjusted $R^{2}$ 0.600 0.600 0.466 0.380 0.282
AIC -6435 -6511 -4018 -4389 -3649
No. of obs. 1147 1147 668 747 628
table 1 aggregates regression results on Eq. (2). Coefficient estimates for expected and unexpected trading volume are positive across all the five exchanges, consistent with the findings in the previous studies on traditional futures market [5, 24].
For exchanges of LOB Model, i.e., Binance and Bybit, all the estimated coefficients measuring the marginal effects of expected and unexpected open interest are negative and significant, while the estimated coefficients for the unexpected portion are also larger in magnitude than those for the expected portion. Since in these exchanges each seller is matched with a buyer of the perpetual futures, aggregated value of long positions equals to that of short positions, so do the open interests. We only include open interests on short in Eq. (2), as open interests on long are of the same value. Since exchanges of LOB Model adopt the limit-order-books, it relates to the determinants of market depth, which Kyle (1985)[13] defines as the volume of unanticipated order flow required to move price by one unit. When capital inflows into the market, the liquidity increases, which is reflected by higher level of open interests, making the price less volatile in response to new orders. Consequently, it is expected that when open interest is large the price volatility conditional on contemporaneous trading volume, which proxies market depth, would be lower. Therefore, those negative estimated coefficients imply that an increase in unexpected open interest, e.g., inflow of speculative capital, lessens the impact of trading volume shock (unexpected volume) on price volatility. In our case, for Binance, when a trader increase his/her position by 1 USD, the marginal effect of the 1 USD increase in trading volume on price volatility is $5.943\times 10^{-8}$, while mitigated by the effect of 1 USD increase in open interest ($-2.365\times 10^{-11}$), resulting in a smaller aggregated effect ($5.943\times 10^{-8}-2.365\times 10^{-11}$). On the one hand, estimated coefficients for trading volumes and open interests have opposite signs, while on the other hand, whether the marginal effect of volume on price volatility is enlarged or mitigated depends on whether open interest is reduced or increased. Trades closing, decreasing, or liquidating positions actually reduce open interest, enlarging the effect of trading volume on price volatility  [^10] Trades opening positions or rising positions increases open interests.. Price would move with larger distance, as less liquidity reacts to absorb a trade. This is reflected by that seven out of eight coefficients for liquidated volume in Binance and Bybit are significant and positive, with expected liquidated volume on short positions for Bybit as the exception. Thus, all else being equal, an unexpected rise in liquidation reduces market depth, via lowering open interest and liquidity.
3.2 Empirical results and analysis for VAMM Model (Perpetual Protocol V2).
Different from the case of LOB Model, where traders are matched by the market maker and provide liquidity, traders in exchanges of VAMM Model are matched with liquidity pools, in which liquidity providers act passively as the counterparty. Traders’ open interest thus does not necessarily influence the market depth in the same way as in LOB Model and the theory of market depth by Kyle (1985)[13] for traditional futures market is inapplicable in this context. For Perpetual Protocol V2, which adopts VAMM as the pricing mechanism, liquidity providers deposit stable coins in the vault based on which the clearing house mints virtual tokens and places those virtual tokens into the Uniswap v3 AMM  [^11] Details of how VAMM works in Perpetual Protocol V2 can be found in https://support.perp.com/hc/en-us/articles/9594157347481-How-It-Works.. For illustration, assume the price of the Bitcoin perpetual is determined by the exchange function:
$Q_{vUSDC}\times Q_{vBTC}=k$ (4)
where $Q_{vUSDC}$ and $Q_{vBTC}$ denote the quantities of virtual USDC (vUSDC) and virtual Bitcoin (vBTC) in the liquidity pool, and $k$ reflects the depth of liquidity, which is determined by liquidity providers. With constant $k$, the price of virtual Bitcoin is determined by the ratio of $Q_{vUSDC}$ and $Q_{vBTC}$:
$P_{vBTC}=\frac{Q_{vUSDC}}{Q_{vBTC}}=k\times Q^{-2}_{vBTC}.$ (5)
The absolute value of the first derivative of $P_{vBTC}$ with respect to $Q_{vBTC}$ is monotonously decreasing, which can be derived as:
$\left|\frac{\partial P_{vBTC}}{\partial Q_{vBTC}}\right|=\left|-2\times k\times Q^{-3}_{vBTC}\right|=2\times k\times Q^{-3}_{vBTC},$ (6)
indicating the decreasing rate of change in pricing as vBTC goes abundant in the liquidity pool. The more vBTC there are in the liquidity pool, the lower volatile its price is. When a trader opens a long position, the clearing house swaps vUSDC for vBTC from the liquidity pool, adding vUSDC to and withdrawing vBTC from the pool. $Q_{vBTC}$ decreases while $P_{vBTC}$ becomes more volatile.
Besides, Uniswap v3 AMM implements concentrated liquidity to enhance capital efficiency, where liquidity providers specify a price range, $[\underline{P},\bar{P}]$, within which to add liquidity  [^12] For details of geometric explanation on concentrated liquidity, readers can refer to [17].. If there is only a single liquidity provider specifying the price interval $\left[P_{vBTC}^{A},P_{vBTC}^{B}\right]$ to add liquidity, where $P_{vBTC}^{A}=\frac{Q_{vUSDC}^{A}}{Q_{vBTC}^{A}}$ and $P_{vBTC}^{B}=\frac{Q_{vUSDC}^{B}}{Q_{vBTC}^{B}}$, this concentrated liquidity can be characterized by drawing a new set of axes, i.e., $\widehat{Q_{vBTC}}$ and $\widehat{Q_{vUSDC}}$ with the origin $\hat{O}$ as shown in Fig. 1, to track the real reserve offered in range $AB$. While the exchange function is defined in the $Q_{vBTC}-Q_{vUSDC}$ space ($virtual$ reserve termed by Unisawp v3), the actual reserve specified in the price range is called $real$. Any point $(Q_{vBTC},Q_{vUSDC})$ in the $virtual$ space can be transformed to a point $(\widehat{Q_{vBTC}},\widehat{Q_{vUSDC}})=(Q_{vBTC}-Q_{vBTC}^{B},Q_{vUSDC}-Q_{vUSDC}^{A})$ in the $real$ space. The liquidity provider actually defines the coordinate of the $real$ origin $\hat{O}$, which is a function of the interval $\left[P_{vBTC}^{A},P_{vBTC}^{B}\right]$.
[…]
Figure 1: A Uniswap v3 pool with a single liquidity provider.
In practice, liquidity offered by multiple liquidity providers cumulates around the current price level, as illustrated in Fig. 2. Trades at the tails of the distribution lead to larger price move, i.e., thus larger price volatility, due to lower liquidity than at the center. Therefore, while the design of concentrated liquidity determines market depth locally, the exchange function Eq. (4) forms price globally. In the long-run, $(Q_{vBTC},Q_{vUSDC})$ moves along the exchange function and the change in slope determines the price volatility. In the short-run, the liquidity distribution centered at the current average price determines price volatility.
Figure 2: Example of Liquidity Distribution on Uniswap v3 AMM. This figure shows the liquidity distribution for the pool of the pair WBTC/USDC on Uniswap V3. This snapshot is made on 7 Nov 2023. Source: https://dune.com/queries/65034/129883.
The implication of Uniswap v3 AMM on pricing volatility is consistent with our empirical findings. In Table 1, the coefficient for expected open interest on short is estimated to be significantly negative ($-1.678\times 10^{-9}$) while that on long is estimated to be significantly positive ($3.569\times 10^{-9}$), indicating a negative association between expected open interest on short and price volatility while a positive association for long. Since the expected portion of open interest reflects traders’ long-run behavior, e.g., hedgers, in cope with the trend, the change could be mainly be described by Eq. (4). When the expected open interest on short rises, the clearing house retrieve vBTC from traders’ account and credit to the balance of the liquidity pool. The increased vBTC balance in the liquidity pool helps decrease the rate of pricing change and thus reducing price volatility according to Eq. (6). Besides, when expected open interest on long increases, the clearing house withdraw vBTC from the pool and credit to traders’, decreasing vBTC balance in the liquidity pool and thus increasing price volatility.
While the expected portion of open interest captures how market status moves along the exchange function specified by Eq. (4), the unexpected portion reflects traders’ unanticipated behavior, e.g., speculation, on the basis of liquidity distribution. In the short-run, liquidity providers are lagged to adjust their bound of price within which to add liquidity, i.e., $[\underline{P},\bar{P}]$, especially when facing unanticipated change. Therefore, open interest on both long and short positions would drive price to the tails of the liquidity distribution, enlarging pricing volatility due to lower liquidity. In Table 1, the estimated coefficients for the association between price volatility and unexpected open interest are all positive, i.e., $1.125\times 10^{-8}$ for short and $1.481\times 10^{-9}$ for long, substantiating our prediction based on the Uniswap v3 AMM pricing mechanism.
In the context of Perpetual Protocol V2, liquidity providers have the capability to contribute liquidity utilizing leverage, thereby incurring a risk of passive liquidation when counterparties are liquidated. Transactions that liquidate liquidity providers’ positions during a price increase are categorized as liquidation on short, whereas those during a price decline are classified as liquidation on long. The liquidation of liquidity providers’ positions results in a reduction of market depth, which is hypothesized to have a direct correlation with increased price volatility and the frequency of liquidations among liquidity providers.
Regarding the expected liquidation volume of traders’ positions, its impact on price volatility is anticipated to be inversely proportional to the effect of increasing expected open interest. Specifically, an expected increase in liquidation on short positions, which reduces open interest in these positions, is posited to have a positive correlation with price volatility, and the converse holds true for long positions. In the case of the unexpected component of liquidation, its influence on price volatility aligns with that of unexpected open interest, with all unforeseen trading activities driving price movements towards the extremities of the liquidity distribution, thereby exacerbating price volatility.
Overall, an increase in unexpected liquidation in both long and short positions is expected to elevate price volatility. This is empirically corroborated in Table 1, where the estimated coefficient for unexpected liquidated volume on long positions is significantly positive ($5.008\times 10^{-7}$). Conversely, the aggregated effect of an increase in expected liquidation on short positions should theoretically be positive. However, this is contradicted by the negative sign of the estimated coefficient for expected liquidated volume on short positions ($-5.983\times 10^{-7}$). A plausible explanation for this anomaly is that a price rise, typically accompanying an increase in short position liquidations, may attract additional liquidity providers. This influx of liquidity providers potentially mitigates price volatility, with their dampening effect on volatility surpassing the amplifying impact of liquidations. We have conducted the Granger Causality Test to examine the relationship between between the return on Bitcoin and the net change in liquidity in the vBTC-vUSDC pool on Perpetual Protocol V2, which you can find in Appendix B. It is shown that return on Bitcoin Granger-cause the net change in liquidity in the vBTC-vUSDC pool, while the reversed relationship does not hold.
The asymmetrical effect of open interest on long and short positions is also predicted in the theoretical work by [3], which suggests that consuming liquidity involves a larger price impact than adding it, resulting in the larger marginal cost of a ETH buy order than the marginal benefit of a ETH sell order. The model proposed in [3] implies an ambiguous reaction of AMM liquidity to asset price volatility, driven by the migration of both traders and liquidity providers between CEXs and DEXs. In our test, since liquidation explicitly includes behavior of both traders and liquidity providers, the ambiguity mentioned in [3] may help explain the inconsistency between our predicted effect of expected liquidation on short positions and the estimated coefficient.
3.3 Empirical results and analysis for Oracle Pricing Model (GMX and GNS).
In exchanges of Oracle Pricing Model such as GMX and GNS, traders accept prices as a given, lacking the capacity to exert direct influence on price movements. The pricing mechanism in these DEXs is governed by a Price Oracle, which assimilates prices from various exchanges, thereby rendering transactions on GMX and GNS incapable of impacting the price directly [^13] The code for GMX’s price feeding smart contract is publicly accessible at: https://github.com/gmx-io/gmx-contracts/blob/master/contracts/oracle/FastPriceFeed.sol.. As a result, trading activities on these platforms, encompassing trading volume and open interest, merely reflect traders’ interpretations and reactions to the information encapsulated within the Price Oracle’s determined prices.
Drawing upon Shalen’s dispersion of beliefs model (1993) [18], which posits a market dichotomy between asymmetrically informed traders, market participants can be categorized into informed and uninformed traders. Uninformed traders, lacking access to private information, often engage in irrational trading based on market noise and tend to overreact to information. They strive, albeit unsuccessfully, to discern private information and market trends from current price fluctuations, leading to heightened position-taking during periods of increased price volatility. In contrast, informed traders, operating on the basis of their private information, exhibit relatively consistent beliefs over time and typically engage in trading within a constrained price range, resulting in a negative correlation between their positions and price volatility [24].
In Table 1, the estimated coefficients that measure the relationship between open interest and price volatility are predominantly negative, except the case of long open interests in GNS. This indicates a more pronounced influence of informed traders’ behavior, leading to the negative signs in these estimations. The impact of uninformed traders on open interest is effectively counterbalanced by the actions of informed traders. Notably, traders’ behaviors exhibit asymmetry between buyers and sellers, as evidenced by less negative (in the case of GMX) or even insignificant (for GNS) coefficients for long positions. This asymmetry could be attributed to uninformed traders’ tendency to overreact more to positive news than negative, leading to an increased propensity for long position accumulation [^14] Existing literature on traditional spot and future markets provides evidence of traders’ asymmetric reactions to positive and negative news [15, 24].. This asymmetry is further evidenced by variations in liquidated volume in relation to price volatility. As indicated in Table 1, for GMX and GNS, the coefficients linking liquidated volume on long positions are significantly positive, whereas those for short positions are not statistically significant.
Aligned with Shalen’s model, it is anticipated that uninformed traders are more likely to increase leverage during periods of heightened price volatility to maximize capital efficiency, thereby enlarging their positions at a lower cost. In contrast, informed traders are less inclined to risk capital under volatile conditions, leading to a negative correlation between leverage and price volatility. Table 2 delineates the estimated results for Eq.(3). All eight coefficients relating price volatility and average leverage are estimated to be negative, indicating a more pronounced change in leverage as influenced by informed traders compared to uninformed traders.
Table 2: Regression results on Eq. (3)
[…]
Intercept [^17] In each cell, the t-statistics are in the parentheses. *, **∗ ∗, and ***∗ ∗ ∗ denote significance at 0.1, 0.05, and 0.01 level, respectively. 2.287e-12 (9.47)*** 0.028 (3.84)*** Lagged volatility 0.229 (7.26)*** 0.369 (10.66)***
Expected portion Unexpected portion
Trading Activity:
Trading volume 1.051e-10 (5.60)*** 5.201e-11 (0.83) 1.359e-10 (10.41)*** 4.593e-10 (6.86)***
OI on short -4.405e-10 (-6.52)*** -1.626e-10 (-1.58) -2.485e-10 (-2.89)*** -4.423e-10 (-2.98)***
OI on long -1.373e-10 (-4.11)*** 3.964e-11 (0.58) -1.091e-10 (-1.36) -5.505e-11 (-0.38)
Liquidation:
Liquidated volume on short positions 2.538e-09 (1.15) -1.467e-09 (-1.27) 1.891e-09 (1.55) 3.177e-10 (0.80)
Liquidated volume on long positions 2.77e-07 (9.14)*** 7.573e-09 (3.45)*** 2.539e-09 (5.06)*** 1.384e-09 (1.42)
Leverage (average level):
Average leverage on short positions -0.0004 (-2.21)** -0.0001 (-1.40) -0.0002 (-3.00)*** -0.0002 (-4.37)***
Average leverage on long positions -0.0004 (-2.15)** -1.193e-05 (-0.36) -0.0003 (-3.65)*** -7.237e-05 (-1.93)**
Adjusted $R^{2}$ 0.398 0.339 AIC -4405 -3693
No. of obs. 747 628
In summary, the analysis reveals that informed traders exhibit a higher level of activity compared to their uninformed counterparts on GMX and GNS, a conclusion substantiated by empirical data pertaining to the dynamics of open interests and leverages. Additionally, the observed asymmetry in open interests and liquidations between long and short positions suggests a tendency among traders to disproportionately overreact to good news. However, it is imperative to acknowledge a critical distinction for analytical purposes: unlike GMX and GNS, centralized exchanges (CEXs) possess a price discovery function, leading to a reciprocal influence between traders’ behavior and market prices. Under Efficient-Market Hypothesis [23], the price aggregate and reflect all the current private information and information contained in the past pricing history. Consequently, it is methodologically unsound to interpret changes in trading activities, e.g., volume, open interest, liquidation, and leverage, solely as reactions to price fluctuations. This perspective challenges the conclusions drawn by [20], who posited that changes in liquidation are merely passive responses to price movements, thereby overlooking the role of liquidation as a contributory factor in price formation within BitMEX, a CEX. This oversight underscores the necessity for a nuanced understanding of the interplay between trader behavior and price dynamics, particularly in the context of CEXs where the price discovery process is inherently more complex.
4 Conclusion
This study underscores the necessity of developing new theoretical models, moving beyond conventional frameworks like Kyle’s model[13], to aptly comprehend the intricacies of DEXs. Our empirical results uncover distinct variations in DEXs, particularly under the VAMM Model, where open interest impacts long and short positions differently. The behavior of traders in exchanges of Oracle Pricing Model (GMX and GNS), who act as price takers, is found to be consistent with Shalen’s dispersion of beliefs model[18]. Our study highlights the significant fundamental difference between CEXs and DEXs, based on which theoretical frameworks specifically for DEX design can be derived in the future works. Table 3 encapsulates a summary of the unique trading behaviors elicited by the varied designs of these models. Any exchange classified into one of our three exchange models is believed to derive distinctive behavior of the traders caused by the design of the exchange model.
Table 3: Summary of Empirical Results
[…]
Analyzed
LOB Model Yes Positively correlated Negatively correlated Symmetrical; Kyle’s model [13] Binance; Bybit
VAMM Model Yes Positively correlated Mixed Asymmetrical; Uniswap v3 AMM Perpetual Prorocol V2
Oracle Pricing Model No Positively correlated Negatively correlated Asymmetrical; Shalen’s dispersion of beliefs model [18] GMX; GNS
5 Acknowledgement
ChatGPT has been used in the writing for the purpose of spelling, syntax and grammar checks. The authors take full responsibility of the output.
[…]